Grade 12th passGeneral Physics

A particle starting from mean position executes SHM with a period 8s .the minimum time in which PE become half of TE

Sneha
8 Years agoGrade 12th pass

1 Answer

Jgun
8 Years ago
POTENTIAL ENERGY + KINETIC ENERGY = TOTAL ENERGY 
TOTAL ENERGY = 2(POTENTIAL ENERGY)
So,POTENTIAL ENERGY = KINETIC ENERGY
So,
mA\omega ^2sin(\omega t)=mA\omega^2cos(\omega t)
\Rightarrow sin(\omega t)=cos(\omega t)
\omega t=\pi/4\Rightarrow(2\pi/8)t=\pi/4\Rightarrow t=1s
 Okay, This is the answer.